There are two investment projects. Both projects will provide benefits to the city for the next 30 years. The annual rate is 7%. Below are the initial cash outlays and monthly benefits:

Ice Cream Shop
Alternative Investment ($K) Monthly Benefits
1 0 0
2 3000 24.5
3 6000 58.5
4 9000 76.4
5 12000 85.5
Cookie Shop
Alternative Investment ($K) Monthly Benefits
1 0 0
2 3000 28.6
3 6000 50.4
4 9000 65.5
5 12000 72.4

a) Find the NPV for each of the alternatives.

b) Based on costs and benefits and considering all possible investments together as one integrated capital program, choose the best alternative under the following total budget constraints. Explain how you arrive at your answers

1) 4 Millon

2) 8 Millon

3) 12 Millon

4) 16 Millon
Answer #1
Ice cream shop
Alternative Investment ($k) Monthly Benefits NPV calculation NPV Ranking Total ranking
1 0 0 0 5
2 3000 24.5 =(PV(7%/12,12*30,-24.5,0))-3000 682.54 4 6
3 6000 58.5 =(PV(7%/12,12*30,-58.5,0))-6000 2,792.99 1 1
4 9000 76.4 =(PV(7%/12,12*30,-76.4,0))-9000 2,483.50 2 2
5 12000 85.5 =(PV(7%/12,12*30,-85.5,0))-12000 851.30 3 5
Cookie shop
Alternative Investment ($k) Monthly Benefits NPV calculation NPV
1 0 0 0 4
2 3000 28.6 =(PV(7%/12,12*30,-28.6,0))-3000 1,298.80 2 4
3 6000 50.4 =(PV(7%/12,12*30,-50.4,0))-6000 1,575.50 1 3
4 9000 65.5 =(PV(7%/12,12*30,-65.5,0))-9000 845.15 3 7
5 12000 72.4 =(PV(7%/12,12*30,-72.4,0))-12000 -1,117.73 5 8
4 Million - Cookie shop alternative 2 as it is the highest available NPV in exchange of 3 million cash outflow
8 Million - Ice cream shop alternative 3 as it is the highest available NPV in exchange of 6 million cash outflow
12 Million - Ice cream shop alternative 3 as it is the highest available NPV in exchange of 6 million cash outflow, and Ice cream alternative 2 and cookie shop alternative 2
16 Million - Ice cream shop alternative 3 as it is the highest available NPV in exchange of 6 million cash outflow, and cookie shop alternative 2 and alternative 3 as it gives highest cumulative NPV